Circular No. 7858

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2002 CW_{134} (LINEAR) An object reported as asteroidal by LINEAR (discovery observation

below), given the designation 2002 CW_{134} on MPS 50314, and later placed on the NEO Confirmation Page, has been found to show cometary activity: Mar. 19.0 UT, diffuse with coma diameter 13′′ (M. Tichý and M. Kočer, Kleť, 1.06-m reflector); Mar. 23.0, diffuse, faint coma extended ∼ 30′′ in p.a. 330^{o} (G. Masi and F. Mallia, Campo Catino, Italy, 0.8-m reflector).

2002 UT  α_{2000}  δ_{2000}  m_{2}
Feb. 7.47271  12^{h}52^{m}46.^{s}90  −5^{o}33′17′′.2  20.1

Astrometry and the following orbital elements appear on MPEC 2002-F49:

T = 2002 Feb. 28.8798 TT  ω = 190.^{o}0804
e = 0.489613  Ω = 348.3172 }2000.0
q = 1.839617 AU  i = 15.2272
a = 3.604353 AU n^{o} = 0.1440334 P = 6.843 years

COMET C/2002 C1 (IKEYA-ZHANG) Ephemeris extension to IAUC 7852 (elements MPEC 2002-F21):

2002 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Mar. 27 1^{h}19.^{m}83 +30^{o}00.^{′}3 0.674 0.542 30.^{o}7 109.^{o}8 3.5
29 1 15.26 +32 43.0 0.643 0.560 31.4 112.0 3.5
31 1 09.42 +35 26.0 0.613 0.582 32.3 113.4 3.6
Apr. 2 1 02.23 +38 08.8 0.586 0.606 33.6 114.0 3.7
4 0 53.59 +40 50.7 0.561 0.632 35.3 113.8 3.8
6 0 43.37 +43 31.2 0.538 0.660 37.4 112.9 3.8
8 0 31.37 +46 09.5 0.517 0.689 39.8 111.5 3.9
10 0 17.36 +48 44.4 0.498 0.720 42.6 109.4 4.1
12 0 01.03 +51 14.4 0.480 0.751 45.8 107.0 4.2
14 23 42.03 +53 37.3 0.464 0.783 49.2 104.1 4.3
16 23 19.97 +55 49.9 0.450 0.816 53.0 100.9 4.4

Visual m_{1} and tail-length estimates: Mar. 7.79 UT, 4.7, 3^{o} (B. H. Granslo, Movatn, Oslo, Norway, 7×50 binoculars); 10.78, 4.4, 5^{o} (M. V. Zanotta, Santa Caterina Valfurva, Italy, 7×42 binoculars); 15.13, 3.7, – (C. S. Morris, Fillmore, CA, naked eye); 19.80, 3.2, 5^{o} (R. J. Bouma, Groningen, The Netherlands, 7×50 binoculars); 21.83, 3.2, 3^{o} (J. J. Gonzalez, Asturias, Spain, 7×50 binoculars).

2002 March 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green